# Vib2Mol: from vibrational spectra to molecular structures—a versatile deep learning model

### v3

@@ Abstract
There will be a paradigm shift in chemical and biological research, to be enabled by autonomous, closed-loop, real-time self-directed decision-making experimentation. Spectrum-to-structure correlation, which is to elucidate molecular structures with spectral information, is the core step in understanding the experimental results and to close the loop. However, current approaches usually divide the task into either database-dependent retrieval and database-independent generation and neglect the inherent complementarity between them. In this study, we proposed Vib2Mol, a versatile deep learning model designed to flexibly handle diverse spectrum-to-structure tasks according to the available prior knowledge by bridging the retrieval and generation. It not only achieves state-of-the-art performance in analyzing theoretical Infrared and Raman spectra, but also outperform previous models at experimental data. Moreover, Vib2Mol demonstrates promising capabilities in predicting reaction products and sequencing peptides, enabling vibrational spectroscopy a real-time guide for autonomous scientific discovery workflows.111codes are available at https://github.com/X1nyuLu/vib2mol
@@ 1 Introduction
With the rapid development of automated experimental design and execution[1, 2], it has become possible to explore potential chemical reactions and study complex life processes with a closed-loop workflow without human intervention. It may significantly accelerate material design and drug discovery. The key to automating such a closed-loop workflow is to design and execute the next experiment on the basis of the prior knowledge. However, this is particularly challenging due to the lack of quantification for the merits of the decisions. In this context, spectra, especially those obtained from in-situ measurements, have become the key to addressing this challenge by providing the basic structural information of molecules thus offering feedback for each decision. Therefore, it is urgent to develop efficient methods to elucidate molecular structures on the basis of spectral information, i.e., spectrum-to-structure correlation.
@@
Leveraging its superior ability to process big data and uncover latent patterns, deep learning (DL) has significantly advanced the spectrum-to-structure tasks. These DL-based methods can be generally categorized into two ways: database–dependent retrieval and database–independent generation. Retrieval-based approaches, including spectrum-spectrum and spectrum-structure retrieval, rely on comparing the to-be-determined spectrum with candidate spectra or molecular structures according to certain rules to find the best match. These approaches are effective in identifying chemicals within the library that has been previously established or delineated on the basis of prior knowledge, such as DeepSearch[3], FastEI[4] and CReSS[5]. However, these methods inevitably face severe limitations when dealing with out-of-library compounds, owing to the big gap between available experimental spectrum-structure pairs ($\sim 10^{6}$) and vast chemical space[6]. In contrast, generation-based approaches, including conditional generation and de novo generation, seek to predict molecular structures directly from spectra, bypassing the establishment and retrieval of databases. These approaches have shown great promise for predicting previously unidentified chemicals[7, 8, 9, 10, 11, 12, 13, 14]. However, the spectral signal obtained from single technique unveils only a partial view of molecular structure. As a result, the process of converting one type of spectral data into its molecular structure is inherently challenging, let alone the complexity and noise in the experimental spectrum.
@@
Indeed, retrieval is efficient enough to determine in-library molecules, whereas generation becomes the only option for interpreting spectra of out-of-library molecules. However, up to now most of the existing methods have either retrieval or generation but not both. Such a paradigm not only makes model unable to provide appropriate solutions as prior knowledge and databases change, but also ignores the synergy between retrieval-based and generation-based spectrum-to-structure tasks, while this synergy could further improve the performance of spectral annotation. As a result, it is ideal to develop a general model that is capable of retrieval and generation simultaneously and provides dynamic solutions on the basis of available knowledge and databases.
@@
In this study, we propose a DL-based vibrational spectrum-to-molecular structure model (Vib2Mol) to flexibly address a variety of spectral annotation tasks according to the available prior knowledge. Vib2Mol adopts an encoder-decoder transformer architecture, and is trained with the strategy of multi-task learning and integrates a wide variety of spectrum-to-structure tasks into one versatile model. For a better evaluation, we compiled theoretical and experimental benchmarks, drawing upon both public datasets and our own Density functional theory (DFT) calculations. Overall, Vib2Mol not only achieves state-of-the-art performance on all benchmarks but also exhibits an overwhelming superiority when compared to mainstream methods. It further enhances the accuracy in interpreting spectra of reaction products and peptide sequencing as more knowledge of target molecules is introduced. This advancement demonstrates significant potential for in-situ intelligent analysis of dynamic chemical transformations and biological processes.
@@ 2 Results
2.1 Multi-task learning framework: correlating vibrational spectrum and molecular structure
@@
The workflow of Vib2Mol during pre-training, including alignment and generation phases, is illustrated in Figure 1. Vib2Mol adopts staged pre-training (SPT) as a fundamental training strategy. In the first stage, the alignment phase (Figure 1A) aims to bring the spectral and structural features of the same molecule as close as possible while separating the features of different molecules simultaneously. Spectra and molecular structures are represented as patch tokens and SMILES tokens, and then encoded into spectral and molecular embeddings by encoders, respectively. These two embeddings are effectively aligned through contrastive learning (CL), enabling cross-modal spectrum-structure retrieval. To further enhance retrieval performance, we deliberately selected hard negative samples—highly similar molecule-spectrum pairs—from each training batch. A matching loss was then utilized to guide the model in learning the subtle distinctions inherent in these challenging samples.
@@
Figure 1B depicts workflow of the second stage, including conditional generation and de novo generation of molecular structures. Conditional generation, i.e., predicting the masked molecular structure on the basis of the spectrum, draws on masked language modeling (MLM). Briefly, SMILES tokens, representing the molecular structure, are randomly masked by 45% and then processed by molecular encoders to generate molecular features. The spectral and molecular encoders were initialized by cloning the parameters from the alignment phase depicted in Figure 1A. These parameters were then subsequently fine-tuned specifically for the generation task to optimize performance. Then molecular decoders fused information from both masked molecular embeddings and spectral features, and predicted the to-be-determined tokens using cross-attention. Differently, de novo generation draws on language modeling (LM). SMILES tokens are sequentially masked from left to right and directly input into the molecular decoders sharing parameters with MLM. Guided by spectral features and previously generated SMILES sequences, the decoders can predict the next SMILES token from left to right until the entire sequence is complete. Note that the chemical formula is an optional input for Vib2Mol, as its provision unequivocally boosts performance. Nevertheless, Vib2Mol remains capable of achieving commendable results even in its absence, which will be discussed in Section 2.3.
@@
Figures 2A to 2E illustrate the workflow of Vib2Mol during application and inference. (1) For spectrum-spectrum retrieval (Figure 2A), instead of directly comparing spectral similarity by metrics such as Pearson correlation coefficient, the to-be-determined spectrum is encoded into an embedding vector, and the cosine similarity is calculated between this vector and the known spectral embedding vector in the database. (2) Spectrum-structure retrieval (Figure 2B) leverages cross-modal retrieval strategy. The spectrum-structure similarity is calculated between the embedding vector of the query spectrum and the molecular embedding vectors of known entities within the database. (3) For conditional generation (Figure 2C), Vib2Mol adopts the encoder-decoder architecture. Both the spectrum and partially masked molecular structure are encoded and then fused through the molecular decoder to generate the SMILES of the masked part. (4) For de novo generation (Figure 2D), Vib2Mol directly employs molecular decoders to sequentially predict each character of SMILES string on the basis of the encoded spectral features until a complete molecular structure is generated. When generating the next token, beam search (see Methods for details) is used to ensure the diversity of the results and then improve the generation performance. A re-ranking module (Figure 2E) is implemented to enhance both spectrum-structure retrieval and de novo generation. Candidate molecules were filtered based on the chemical formula, and then a pre-trained molecular encoder functions as a matching module, generating scores from a comprehensive evaluation of the query spectrum and candidate molecule features. Only candidates exhibiting high matching scores are subsequently selected as the final results. Even in the absence of chemical formula, the re-ranking module can still effectively sort candidates relying solely on its model-based scoring.
@@
The model’s inherent parameter sharing and feature reusing allow it to address these four spectrum-to-structure tasks, without the need for additional fine-tuning or training. No wondering, spectrum-structure retrieval is more effective than spectrum-spectrum retrieval because it makes better use of molecular databases[15]. De novo generation offers more flexibility than conditional generation as it does not require a predefined molecular scaffold to predict side-chain structures. Therefore, for the spectrum-to-structure problem, spectrum-structure retrieval and de novo generation constitute the most versatile solutions, and will be the primary focus of following discussion.
@@ 2.2 State-of-the-art performance of Vib2Mol
To evaluate the performance of Vib2Mol on different spectrum-to-structure tasks and fairly compare it with current advanced models, five benchmarks are established. The benchmarks, which are described in detail in Methods, consist of three theoretical datasets (QM9S[16], VB-Mols, and VB-GEOM) and two experimental ones (SDBS[17] and NIST-IR[18]). Cumulatively, these benchmarks provide a total of 662,668 theoretical and 17,774 experimental spectra. There benchmarks enable a fair and comprehensive comparison with existing state-of-the-art methods, which can be broadly classified into three categories (see Methods for details): retrieval model (vibraCLIP[19]), generative models (PBSA[9], IR2Mol[20]), and comprehensive models capable of both retrieval and generation tasks (SMEN21, Vib2Mol). All benchmarking metrics are illustrated in Figure 2 and meticulously documented in Table S1-S5. Unless otherwise noted, Top-1 Recall (Recall@1) serves as the default metric throughout this section.
@@
The generalization of model was evaluated on the VB-GEOM benchmark, which differs from QM9S and VB-Mols in three ways. (1) As an out-of-distribution dataset, VB-GEOM provides a rigorous test of the robustness of models. (2) A different task, i.e., spectrum-spectrum retrieval is offered. (3) A zero-shot evaluation is performed, which means the model was pre-trained on the VB-Mols dataset and then directly evaluated on the VB-GEOM test set without any fine-tuning on VB-GEOM’s data. Vib2Mol demonstrated superior performance (77.54% for Raman, 75.33% for IR) among all models again, outperforming the second-place vibraCLIP (74.96% for Raman, 64.43% for IR) by a significant margin. In contrast, SMEN performed poorly (41.80% for Raman, 49.01% for IR), barely exceeding traditional cosine similarity and Pearson Correlation Coefficient, which represent the lower bound for this benchmark (Table S3).
@@
When it comes to experimental benchmarks, all models were pre-trained on VB-Mols and then fine-tuned with the corresponding training set, and finally evaluated with the test set, given the inherent data scarcity of these experimental datasets. Vib2Mol demonstrated a substantially expanded lead in these benchmarks. Our model achieved the Recall@1 of 83.54%, 86.17%, and 90.43% on NIST-IR, SDBS-IR, and SDBS-Raman, respectively, which surpassing the metrics of vibraCLIP by almost 40 percentage points. Notably, SMEN was excluded since it could not be fine-tuned without molecular conformation data for target spectra in experimental benchmarks.
Figure 3B illustrates the de novo generation performance of various models across all benchmarks. Vib2Mol consistently achieved state-of-the-art performance. For the QM9S and VB-Mols benchmarks, Vib2Mol outperformed the second-best model by at least 10 percentage points. This performance gap was particularly evident in the VB-GEOM benchmark, which requires models to generate molecules from out-of-library spectra under zero-shot conditions. In this challenging scenario, Vib2Mol achieved a remarkable Recall@1 of 66.86% and 63.07% for Raman and infrared spectra, respectively. In contrast, the PBSA model yielded significantly lower metrics (50.90% and 47.32%), while the IR2Mol and SMEN fail to work on this scenario, resulting in a Recall@1 of 0%.The better performance of our model than others can be attributed to the meticulously designed pre-training strategy of Vib2Mol. Specifically, cross-modal alignment enables Vib2Mol to learn generalizable spectral and structural representations, while multi-task learning prevents overfitting to specific patterns in the training data.
We further examined the impact of integrating multi-modal spectral information on Vib2Mol performance. Figures 3C and 3D compare the retrieval and generation performance of Vib2Mol with multi-modal spectral input (Vib2Mol-MM) against Raman-only (Vib2Mol-Raman) and IR-only (Vib2Mol-IR) inputs. Details of processing multi-modal spectral signals can be found in supplementary information. A consistent trend emerged across all datasets: models based on Raman outperformed those based on IR, while models integrating both modalities significantly surpassed methods relying on single-modal input, which is consistent with related work[21]. Take de novo generation on the SDBS benchmark as an example, Vib2Mol-IR and Vib2Mol-Raman achieved Recall@1 scores of 52.84% and 56.03%, respectively, whereas, Vib2Mol-MM reached 68.44%, clearly demonstrating the benefit of multi-modal integration. The diverse molecular structural information provided by different spectra offers the model richer clues and a more integrated framework for constructing molecular structures or spectra. To further enhance the accuracy of structural elucidation, one could consider incorporating other spectral information from different modalities, such as NMR and MS, to provide more comprehensive perspectives on molecular structure features.

### v4

@@ Abstract
There will be a paradigm shift in chemical and biological research, to be enabled by autonomous, closed-loop, real-time self-directed decision-making experimentation. Spectrum-to-structure correlation, which is to elucidate molecular structures with spectral information, is the core step in understanding the experimental results and to close the loop. However, current approaches usually divide the task into either database-dependent retrieval and database-independent generation and neglect the inherent complementarity between them. In this study, we proposed Vib2Mol, a unified deep learning framework designed to flexibly handle diverse spectrum-to-structure tasks according to the available prior knowledge by bridging the retrieval and generation. Empowered by our coarse-to-fine retrieval and generate-then-rerank strategies, Vib2Mol not only achieves state-of-the-art performance in analyzing theoretical Infrared and Raman spectra, but also outperform previous models on experimental data. Moreover, our model demonstrates promising capabilities in predicting reaction products and sequencing peptides, enabling vibrational spectroscopy a potential guide for autonomous scientific discovery workflows.
@@ 1 Introduction
With the rapid development of automated experimental design and execution[burger2020mobile, dai2024autonomous], it has become possible to explore potential chemical reactions and study complex life processes with a closed-loop workflow without human intervention. It may significantly accelerate material design and drug discovery. The key to automating such a closed-loop workflow is to design and execute the next experiment on the basis of the prior knowledge. However, this is particularly challenging due to the lack of quantification for the merits of the decisions. In this context, spectra, especially those obtained from in-situ measurements, have become the key to addressing this challenge by providing the basic structural information of molecules thus offering feedback for each decision. Therefore, it is urgent to develop efficient methods to elucidate molecular structures on the basis of spectral information, i.e., spectrum-to-structure correlation.
@@
Leveraging its superior ability to process big data and uncover latent patterns, deep learning (DL) has significantly advanced the spectrum-to-structure tasks. These DL-based methods can be generally categorized into two ways: database–dependent retrieval and database–independent generation. Retrieval-based approaches, including spectrum-spectrum and spectrum-structure retrieval, rely on comparing the to-be-determined spectrum with candidate spectra or molecular structures according to certain rules to find the best match. These approaches are effective in identifying chemicals within the library that has been previously established or delineated on the basis of prior knowledge, such as DeepSearch[yu2025towards], FastEI[yang2023ultra] and CReSS[yang2021cross]. However, these methods inevitably face severe limitations when dealing with out-of-library compounds, owing to the big gap between available experimental spectrum-structure pairs ($\sim 10^{6}$) and vast chemical space[reymond2015chemical]. In contrast, generation-based approaches, including conditional generation and de novo generation, seek to predict molecular structures directly from spectra, bypassing the establishment and retrieval of databases. These approaches have shown great promise for predicting previously unidentified chemicals[hu2024accurate, litsa2023end, wu2025transformer, stravs2022msnovelist, tran2017novo, mao2023mitigating, qiao2021computationally, alberts2024leveraging]. However, the spectral signal obtained from single technique unveils only a partial view of molecular structure. As a result, the process of converting one type of spectral data into its molecular structure is inherently challenging, let alone the complexity and noise in the experimental spectrum.
@@
Indeed, retrieval is efficient enough to determine in-library molecules, whereas generation becomes the only option for interpreting spectra of out-of-library molecules. However, up to now most of the existing methods have either retrieval or generation but not both. Such a paradigm not only makes model unable to provide appropriate solutions as prior knowledge and databases change, but also ignores the synergy between retrieval-based and generation-based spectrum-to-structure tasks, while this synergy could further improve the performance of spectral annotation. As a result, it is ideal to develop a unified framework capable of simultaneous retrieval and generation, providing dynamic solutions on the basis of available knowledge and databases while ensuring minimal overhead in parameter volume and training costs.
@@
In this study, we propose a DL-based vibrational spectrum-to-molecular structure framework (Vib2Mol) to flexibly address a variety of spectral annotation tasks according to the available prior knowledge. Built upon an encoder-decoder Transformer architecture, Vib2Mol utilizes a multiphase training consisting of alignment and generation phases. By employing multi-task learning at each phase, the framework effectively unifies diverse spectrum-to-structure tasks into one comprehensive model. In addition, we implement coarse-to-fine retrieval and generate-then-rerank strategies to further enhance spectrum-structure retrieval and de novo generation, respectively. For a better evaluation, we compiled theoretical and experimental benchmarks, drawing upon both public datasets and our own Density functional theory (DFT) calculations. Overall, Vib2Mol not only achieves state-of-the-art performance on all benchmarks but also exhibits an overwhelming superiority when compared to mainstream methods. It further enhances the accuracy in interpreting spectra of reaction products and peptide sequencing as more knowledge of target molecules is introduced. This advancement demonstrates significant potential for in-situ intelligent analysis of dynamic chemical transformations and biological processes.
@@ 2 Results
2.1 Workflow of Vib2Mol
@@
The workflow of Vib2Mol during training, including alignment and generation phases, is illustrated in Figure 1. Vib2Mol adopts multiphase training (MPT) and employs multi-task learning during each phase. The first alignment (Figure 1A) aims to bring the spectral and structural features of the same molecule as close as possible while separating the features of different molecules simultaneously. Spectra and molecular structures are represented as patch tokens and SMILES tokens, and then encoded into spectral and molecular embeddings by encoders, respectively. These two embeddings are effectively aligned through contrastive learning (CL), enabling cross-modal spectrum-structure retrieval. To further enhance retrieval performance, we deliberately selected hard negative samples—highly similar molecule-spectrum pairs—from each training batch. A matching loss was then utilized to guide the model in learning the subtle distinctions inherent in these challenging samples.
@@
Figure 1B depicts workflow of the second phase, including conditional generation and de novo generation of molecular structures. Conditional generation, i.e., predicting the masked molecular structure on the basis of the spectrum, draws on masked language modeling (MLM). Briefly, SMILES tokens, representing the molecular structure, are randomly masked by 45% and then processed by molecular encoders to generate molecular features. The spectral and molecular encoders were initialized by cloning the parameters from the alignment phase depicted in Figure 1A. These parameters were then subsequently fine-tuned specifically for the generation task to optimize performance. Then molecular decoders fused information from both masked molecular embeddings and spectral features, and predicted the to-be-determined tokens using cross-attention. Differently, de novo generation draws on language modeling (LM). SMILES tokens are sequentially masked from left to right and directly input into the molecular decoders sharing parameters with MLM. Guided by spectral features and previously generated SMILES sequences, the decoders can predict the next SMILES token from left to right until the entire sequence is complete. Note that the chemical formula is an optional input, as vibrational spectroscopies like Raman and IR are inherently limited in determining molecular formulas. While including the formula unequivocally enhances performance, Vib2Mol remains capable of achieving commendable results in its absence (Section 2.4).
@@
Figures 2A to 2E illustrate the workflow of Vib2Mol during inference. (1) For spectrum-spectrum retrieval (Figure 2A), instead of directly comparing spectral similarity by metrics such as Pearson correlation coefficient, the to-be-determined spectrum is encoded into an embedding vector, and the cosine similarity is calculated between this vector and the known spectral embedding vector in the database. (2) Spectrum-structure retrieval (Figure 2B) leverages coarse-to-fine retrieval strategy. Candidate molecules are identified by calculating the similarity between the embedding vector of the query spectrum and those of known molecules within the database (coarse retrieval). Subsequently, the matching module acts as a reranker to identify the most accurate structure from the candidate pool (fine-grained reranking). (3) For conditional generation (Figure 2C), Vib2Mol adopts the encoder-decoder architecture. Both the spectrum and partially masked molecular structure are encoded and then fused through the molecular decoder to generate the SMILES of the masked part. (4) For de novo generation (Figure 2D), Vib2Mol follows a generate-then-rerank strategy. In the generation step, the molecular decoder sequentially predicts each character of the SMILES string on the basis of the encoded spectral features. During token generation, beam search (see Methods for details) is employed to maintain output diversity and enhance generative performance. Subsequently, consistent with the pipeline of spectrum-structure retrieval, the matching module acts as a reranker to identify the most accurate structure from the candidate pool. Figure 2E illustrated the implementation of reranker. Candidate molecules were filtered with their chemical formula, and then a pretrained molecular encoder functions as a matching module, generating scores from a comprehensive evaluation of the query spectrum and candidate molecule features. Only candidates exhibiting high matching scores are subsequently selected as the final results. Even in the absence of chemical formula, the reranking module can still effectively sort candidates relying solely on its model-based scoring.
@@
TIt is noteworthy that extensive parameter sharing and feature reuse are employed during both training and inference. Consequently, Vib2Mol not only operates with less active parameters (see Methods for details) but also effectively addresses these four spectrum-to-structure tasks. Among these, spectrum-structure retrieval is more effective than spectrum-spectrum retrieval because it makes better use of molecular databases[lu2024deep]. De novo generation offers more flexibility than conditional generation as it does not require a predefined molecular scaffold to predict side-chain structures. Therefore, for the spectrum-to-structure problem, spectrum-structure retrieval and de novo generation constitute the most versatile solutions, and will be the primary focus of following discussion.
@@ 2.2 State-of-the-art performance of Vib2Mol
To evaluate the performance of Vib2Mol on different spectrum-to-structure tasks and fairly compare it with current advanced models, five benchmarks are established. The benchmarks, which are described in detail in Methods, consist of three theoretical datasets (QM9S[zou2023deep], VB-Mols, and VB-GEOM) and two experimental ones (SDBS[saito2011development] and NIST-IR[linstorm1998nist]). Cumulatively, these benchmarks provide a total of 662,668 theoretical and 17,774 experimental spectra. There benchmarks enable a fair and comprehensive comparison with existing state-of-the-art methods, which can be broadly classified into three categories (see Methods for details): retrieval model (vibraCLIP[rocabert2025multi]), generative models (PBSA[wu2025transformer], IR2Mol[alberts2025setting]), and comprehensive models capable of both retrieval and generation tasks (SMEN21, Vib2Mol). The hyperparameters of these models are recorded in Table S1-S2. All benchmarking metrics are illustrated in Figure 2 and meticulously documented in Table S3-S5. Unless otherwise noted, Top-1 Recall (Recall@1) serves as the default metric throughout this section.
@@
The spectrum-spectrum retrieval performance was further evaluated using VB-GEOM, a subset derived from the VB-Mols test set. This benchmark consists of pairs of stable conformers for each molecule, and each individual conformer is linked to a specific IR/Raman spectrum. The task is how to accurately identify the specific reference matching the query spectrum while distinguishing the query spectrum from structurally or spectrally similar incorrect matches. Vib2Mol demonstrated superior performance (77.54% for Raman, 75.33% for IR) among all models again, outperforming the second-place vibraCLIP (75.22% for Raman, 70.16% for IR) by a significant margin (Table S5).
@@
When it comes to experimental benchmarks, all models were pretrained on VB-Mols and then fine-tuned with the corresponding training set, and finally evaluated with the test set, given the inherent data scarcity of these experimental datasets. Vib2Mol demonstrated a substantially lead in these benchmarks. Our model achieved the Recall@1 of 83.54%, 86.17%, and 90.43% on NIST-IR, SDBS-IR, and SDBS-Raman, respectively, which surpassing the metrics of vibraCLIP. Notably, SMEN was excluded since it could not be fine-tuned without molecular conformation data for target spectra in experimental benchmarks.
Figure 3B illustrates the de novo generation performance of various models across all benchmarks. Vib2Mol consistently achieved state-of-the-art performance. For the QM9S and VB-Mols benchmarks, Vib2Mol outperformed the second-best model by at least 10%. The distinguished performance of Vib2Mol can be attributed to not only multi-task learning and the multiphase training but also the generate-then-rerank strategy during inference.
However, we observed a pronounced performance gap between theoretical and experimental spectral test sets, owing to the following three factors: 1. Critical role of data quality and consistency. Different from the pristine mapping between structure and spectrum in the theoretical spectral dataset, experimental spectra (such as SDBS and NIST) are crowd-sourced and inherently heterogeneous, considering the variance in experimental conditions (e.g., energy resolution, quantum efficiency of CCD, temperature and solvent) can obscure the underlying spectrum-structure correlation; 2. Sparsity and incompleteness of experimental spectra. Experimental spectra frequently fail to provide full-range spectral coverage, unlike their theoretical counterparts; 3. Data scarcity in the experimental domain. For instance, sample sizes are several orders of magnitude smaller than theoretical datasets, which limits the ability of model to generalize complex patterns from theoretical simulations to the noisy, low-resource realities of experimental measurement, potentially leading to overfitting.
2.3 Unified multimodal modeling and large-scale retrieval
The impact of multimodal spectral information on the performance of Vib2Mol was further investigated. Rather than maintaining separate models for Raman-only, IR-only, or joint inputs, we developed a unified, multimodal model capable of intelligently processing any of these scenarios. This was achieved by employing masked modality modeling (MMM) during the training of multimodal variant of Vib2Mol (Vib2Mol-MM). Specifically, in 50% of the training steps, Vib2Mol-MM was provided with joint IR and Raman signals. For the remaining half, only one modality was randomly selected as input (see Methods for implementation details). This strategy not only enables the model to handle diverse input combinations seamlessly, but also reduces the cost of model storage and maintenance by shifting the spatial complexity from $O(2^{M}-1)$ to constant $O(1)$ where M denotes the number of spectral modalities.
Figures 3C and 3D compare the retrieval and generation performance of Vib2Mol-MM with and without MMM. A trend consistent with related work[hu2025deep] emerged across all datasets: Raman-based models outperformed IR-based spectra, and multimodal approach (IR+Raman) substantially outperformed unimodal models across all evaluated benchmarks. Obviously, the complementary molecular structural information provided by these two vibrational spectra offers the model richer clues and a more integrated framework for constructing molecular structures or spectra. In addition, the impact of MMM on performance is significant, particularly for Raman-only and IR-only tasks on experimental data. Without MMM, a Raman-based model is restricted to learning spectrum-structure correlations solely from Raman data. In contrast, the introduction of MMM enables Vib2Mol-MM to learn a generalized representation of these correlations by leveraging the underlying physical information from both Raman and IR modalities, thereby enhancing the model’s robustness on single-modality tasks.
[…]